# THE MORE, THE MERRIER: CONTRASTIVE FUSION FOR HIGHER-ORDER MULTIMODAL ALIGNMENT

### v1

@@ 5.1 AV-MNIST
MOSI |
Trimodal CLIP | 22.9 1.9 | 20.5 1.6 | – | 23.5 3.6 | 23.3 1.8 | – | 20.9 2.0 | 24.3 2.1 | – | 22.6 | – |
| Symile [saporta2024contrasting] | - | - | 16.3 0.9 | - | - | 18.1 2.4 | - | - | 17.1 1.7 | – | 17.17 | |
| GRAM [cicchetti2024gramian] | 6.0 3.1 | 10.3 2.9 | 16.3 2.5 | 7.7 2.3 | 0.3 0.6 | 7.9 3.4 | 12.2 4.1 | 0.2 0.2 | 12.0 4.2 | 6.1 | 10.1 | |
| TRIANGLE [cicchetti2025triangle] | - | - | 8.3 1.3 | - | - | 4.9 0.7 | - | - | 8.8 1.4 | – | 7.33 | |
@@
UR-FUNNY |
Trimodal CLIP | 3.7 0.6 | 4.0 0.6 | – | 3.8 0.5 | 16.2 1.6 | – | 4.0 0.4 | 16.8 1.2 | – | 6.5 | – |
| Symile [saporta2024contrasting] | - | - | 3.7 0.5 | - | - | 15.4 0.9 | - | - | 16.5 0.6 | – | 11.87 | |
| GRAM [cicchetti2024gramian] | 3.2 0.5 | 3.0 0.9 | 3.9 0.8 | 3.2 0.9 | 0.1 0.2 | 3.1 0.4 | 3.9 0.4 | 0.1 0.0 | 3.3 0.5 | 2.8 | 3.0 | |
| TRIANGLE [cicchetti2025triangle] | - | - | 4.2 0.4 | - | - | 3.3 0.7 | - | - | 3.6 0.9 | – | 3.7 | |
@@
MUStARD |
Trimodal CLIP | 70.7 4.1 | 29.1 3.6 | – | 70.5 4.3 | 26.5 3.2 | – | 30.4 3.4 | 24.6 2.8 | – | 42.0 | – |
| Symile [saporta2024contrasting] | - | - | 61.5 6.3 | - | - | 57.0 3.0 | - | - | 21.3 4.4 | – | 46.6 | |
| GRAM [cicchetti2024gramian] | 61.7 8.0 | 24.8 5.3 | 81.0 4.1 | 67.2 5.4 | 24.8 5.3 | 67.8 5.4 | 27.9 8.7 | 5.1 2.1 | 31.2 9.2 | 35.2 | 47.9 | |
| TRIANGLE [cicchetti2025triangle] | - | - | 68.7 4.2 | - | - | 59.9 3.7 | - | - | 21.9 5.0 | – | 29.4 | |
@@ 5.2 Affective Computing Benchmarks
We next evaluate on the MOSI, UR-FUNNY, and MUStARD datasets, which involve text, audio, and video for multimodal affect understanding. We assess (i) 11 retrieval, (ii) 21 retrieval, and (iii) multimodal classification via a linear probe.
As shown in Table 2, ConFu consistently improves cross-modal retrieval performance. In particular, the 21 retrieval setting surpasses 11 retrieval on UR-FUNNY and MUStARD, suggesting that combining two input modalities provides richer and more discriminative cues. These results indicate that ConFu effectively exploits complementary multimodal information rather than relying on redundant correlations, leading to stronger alignment and more stable performance across runs.
| Method | MOSI | UR-FUNNY | MUStARD |
| Tri-CLIP | 63.5 2.1 | 64.0 0.9 | 62.1 3.8 |
| Symile [saporta2024contrasting] | 67.5 1.3 | 64.7 1.0 | 60.5 5.0 |
| GRAM [cicchetti2024gramian] | 65.5 2.5 | 64.8 0.9 | 64.7 2.3 |
| TRIANGLE [cicchetti2025triangle] | 65.0 2.8 | 64.5 0.4 | 64.6 2.5 |
| ConFu | 66.7 2.1 | 64.9 1.0 | 64.1 2.5 |
Notably, GRAM, the only baseline supporting both retrieval configurations, struggles with 11 alignments in general and fails completely for modalities M2 and M3 on MOSI and UR-FUNNY suggesting that balancing 21 and 11 alignment is non-trivial. Furthermore, methods such as TRIANGLE and Symile, which focus on trimodal alignment without explicitly modeling pairwise relations, perform consistently worse than ConFu. Overall, ConFu ranks first or second across nearly all querytarget retrieval settings, in both 11 and 21 configurations, demonstrating robust and consistent cross-modal alignment. Moreover, Table 3 reports linear-probe classification results, showing that ConFu attains the highest accuracy on UR-FUNNY and remains competitive on MUStARD and MOSI. This demonstrates that fused supervision benefits retrieval while also improving the quality of multimodal representations.
@@ 5.3 Fine-Grained Bird Classification
| SSW60 Acc. (%) | VB100 Acc. (%) | |||||
| Method | A | V | A+V | A | V | A+V |
| CLIP | 29.9 | 70.1 | – | 4.2 | 20.6 | – |
| Tri-CLIP | 31.1 | 69.0 | – | 3.9 | 20.7 | – |
| Symile [saporta2024contrasting] | - | - | 60.2 | - | - | 13.4 |
| TRIANGLE [cicchetti2025triangle] | - | - | 64.1 | - | - | 12.1 |
| GRAM [cicchetti2024gramian] | 0.7 | 66.6 | 56.9 | 1.3 | 13.7 | 8.0 |
| ConFu | 30.3 | 69.4 | 71.44 | 3.4 | 19.3 | 18.1 |
ConFu is further evaluated on fine-grained bird classification using the VB100 and SSW60 datasets. Models are pretrained on Bird-MML, our curated dataset of 150K audio–image–text triplets and assessed under zero-shot and few-shot regimes. We consider two evaluation settings: (1) the multi-frame protocol from [van2022exploring], where each video clip is divided into eight segments, the middle frame of each segment is sampled, and the mean embedding of the eight frames is used as the visual representation; and (2) a single-frame setting, providing a more constrained visual input per video. Table 4 reports zero-shot results in the multi-frame setting. ConFu achieves the highest accuracy on SSW60 (71.44%), confirming that multimodal fusion improves fine-grained discrimination when both modalities are informative. On VB100, where the audio modality is largely uninformative ( unimodal accuracy), the fused model (A+V) maintains performance close to the best visual-only baseline (18.15% vs. 20.69%), demonstrating robustness to distracting modalities. In contrast, GRAM, Symile, and TRIANGLE exhibit a notable decline in accuracy (5–10%) relative to ConFu. Few-shot trends in Fig. 4 mirror these findings. Fusion offers clear advantages on SSW60 and stable performance on VB100, showing that ConFu generalizes well even when one modality is weak.
In the single-frame setting, designed to evaluate performance with limited visual information, we averaged classification results across 19 uniformly sampled frames to reduce bias. As shown in Table 5, the fused (A+V) model exceeds the best baseline by roughly 4% on SSW60 and complementary audio–visual cues when visual input is sparse. On VB100, performance drops slightly (by 2–3%) due to the weaker audio modality ( accuracy). Overall, ConFu benefits from richer multimodal supervision, maintaining strong robustness under modality degradation while preserving scalability and simplicity. Additional results on a few-shot adaptation can be found in the Appendix C.1.
@@
| Symile [saporta2024contrasting] | - | - | 61.4 0.2 | - | - | 13.2 0.4 |
| TRIANGLE [cicchetti2025triangle] | - | - | 58.6 3.3 | - | - | 8.7 0.4 |
| GRAM [cicchetti2024gramian] | 0.7 | 59.0 3.3 | 54.9 3.2 | 1.3 | 12.9 0.4 | 7.6 0.3 |
@@ 6 Ablation
6.1 Modality competition
Although multimodal learning seeks to combine complementary cues, models often show modality dominance, where stronger signals suppress weaker ones [huang2022modality]. To assess how our fusion is affected by this effect, we analyze in Fig. 5 the overlap of prediction errors across modalities and fusion. Results reveal visual dominance: The larger portion of correct samples (39%) are predicted by both the fusion and vision models, while the datapoints that only audio predicts correctly or fusion-audio overlap remain small (5–6%), and purely audio-driven successes are rare (2%). Nevertheless, about 5% of samples are uniquely solved by the audiovisual model, indicating that fusion captures complementary information beyond either modality alone. Finally, some classes benefit substantially from fusion (e.g., Label 31 see Appendix B.5), while others remain uniformly difficult.
6.2 Ablation on Noise-Induced Distribution Shift
@@
| Method | Mod. | V noise |
| Tri-CLIP (V) | V | 35.4 0.4 |
| Symile [saporta2024contrasting] | A+V | 40.9 0.3 |
| GRAM [cicchetti2024gramian] | A+V | 25.9 0.1 |
| TRIANGLE [cicchetti2025triangle] | A+V | 26.5 0.2 |
| ConFu | A+V | 45.4 0.1 |
@@ 7 Discussion
Empirical results across synthetic and real-world datasets demonstrate that ConFu effectively captures complementary cross-modal structure. On the synthetic XOR benchmark, ConFu successfully models the underlying multimodal dependencies that pairwise and geometric baselines miss even with higher-dimensional representations (Appendix B.2). In AV-MNIST, ConFu dominates unimodal approaches in zero-shot classification, indicating successful utilization of both modalities. The 21 retrieval evaluations on UR-FUNNY, and MUStARD further confirm that combining two query modalities provides richer semantic grounding than either considered in isolation.
An interesting observation in the 21 case on the MOSI dataset is that fusion slightly underperforms compared to the pairwise baseline. Remarkably, this issue is substantially alleviated by introducing partial modality masking (Appendix B.3), supporting the hypothesis that suboptimal performance stems from shortcut learning in cases where redundant information dominates.
@@
Limitations and Future Work.
ConFu is well-positioned to scale to any number of modalities, offering flexibility for diverse multimodal tasks. However, as the number of modalities increases, computational demands may also rise. Depending on the task, selectively pruning certain alignment terms, shown in Appendix E, can help maintain efficiency. For example, omitting many-to-many alignments, when assuming that only single-modality queries or targets will be needed, opens a promising avenue for future adaptation and research.
ConFu currently relies on fully aligned modalities during training, and exploring ways to relax or bypass this requirement remains crucial for many applications. In particular, enabling ConFu to handle partially missing modalities, such as training with incomplete triplets or drawing on ideas related to pseudopair construction [wang2024omnibind, zhang2024extending, wang2023connecting], could potentially make the framework more robust and broadly applicable across real-world multimodal settings.
Further prospective directions could include extending ConFu to incorporate factorized representations inspired by recent factorization-based and redundancy-reduction approaches [liang2023factorized, wang2024decoupling, dufumier2024align], developing adaptive mechanisms to regulate modality competition [kontrasbalancing, MLB, OGMGE, ReconBoost, wei2024diagnosing], and investigating how ConFu can enhance robustness in multimodal retrieval, particularly when one or more modalities are noisy, incomplete, or severely corrupted.
@@ A.2 MOSI, MUStARD, and UR-FUNNY
For these datasets, all models were trained for 100 epochs, producing 256-dimensional representations. We used a batch size of 128 and a learning rate of . Following prior work [liang2023factorized], each modality encoder was implemented as a 5-layer Transformer with 5 attention heads.
@@ B.1 Effect of the Parameter
The results in Table 7 highlight the influence of the parameter, which balances the contribution of modality-specific and shared representations. These values were selected from the Pareto front obtained by jointly optimizing the two retrieval objectives—mean Recall@10 for 11 and 21 tasks. We observe that moderate values (0.1–0.5) yield the best trade-off between the two objectives, and that the optimal varies across datasets, indicating optimal depends on data distribution.
@@ B.4 Noise experiments
| Symile [saporta2024contrasting] | 58.4 | 21.2 | 59.5 | 27.9 | 60.0 | 40.9 |
| GRAM [cicchetti2024gramian] | 58.9 | 4.0 | 58.3 | 8.94 | 58.1 | 25.9 |
| TRIANGLE [cicchetti2025triangle] | 63.9 | 3.4 | 64.0 | 7.24 | 64.0 | 26.5 |
@@ C.1 Few shot adaptation on SSW60, VB100, CUB200 complete results
In this section, we present the complete results of our few-shot adaptation experiments on the SSW60, VB100, and CUB200 [wah2011caltech] datasets. The following figures include all baseline methods for comparison and illustrate adaptation performance across varying number of shot settings. These results provide a comprehensive view of how each method generalizes to limited-data scenarios and adapts to novel classes.
@@ D.2 Affective Computing Benchmarks
MultiBench [liang2021multibench] is a collection of benchmarking datasets designed to evaluate multimodal representation learning across a wide range of modalities and domains. In our experiments, we used the MOSI, UR-FUNNY, and MUStARD datasets. MultiBench provides pre-extracted features and predefined dataset splits, both of which we adopt in our setup.

### v2

@@ 5.1 AV-MNIST
| MOSI | Trimodal CLIP | 22.9 1.9 | 20.5 1.6 | – | 23.5 3.6 | 23.3 1.8 | – | 20.9 2.0 | 24.3 2.1 | – | 22.6 | – |
| Symile [27] | - | - | 16.3 0.9 | - | - | 18.1 2.4 | - | - | 17.1 1.7 | – | 17.2 | |
| GRAM [6] | 6.0 3.1 | 10.3 2.9 | 16.3 2.5 | 7.7 2.3 | 0.3 0.6 | 7.9 3.4 | 12.2 4.1 | 0.2 0.2 | 12.0 4.2 | 6.1 | 12.1 | |
| TRIANGLE [5] | - | - | 8.3 1.3 | - | - | 4.9 0.7 | - | - | 8.8 1.4 | – | 7.3 | |
@@
| UR-FUNNY | Trimodal CLIP | 3.7 0.6 | 4.0 0.6 | – | 3.8 0.5 | 16.2 1.6 | – | 4.0 0.4 | 16.8 1.2 | – | 8.1 | – |
| Symile [27] | - | - | 3.7 0.5 | - | - | 15.4 0.9 | - | - | 16.5 0.6 | – | 11.9 | |
| GRAM [6] | 3.2 0.5 | 3.0 0.9 | 3.9 0.8 | 3.2 0.9 | 0.1 0.2 | 3.1 0.4 | 3.9 0.4 | 0.1 0.0 | 3.3 0.5 | 2.3 | 3.4 | |
| TRIANGLE [5] | - | - | 4.2 0.4 | - | - | 3.3 0.7 | - | - | 3.6 0.9 | – | 3.7 | |
@@
| MUStARD | Trimodal CLIP | 70.7 4.1 | 29.1 3.6 | – | 70.5 4.3 | 26.5 3.2 | – | 30.4 3.4 | 24.6 2.8 | – | 42.0 | – |
| Symile [27] | - | - | 61.5 6.3 | - | - | 57.0 3.0 | - | - | 21.3 4.4 | – | 46.6 | |
| GRAM [6] | 61.7 8.0 | 24.8 5.3 | 81.0 4.1 | 67.2 5.4 | 24.8 5.3 | 67.8 5.4 | 27.9 8.7 | 5.1 2.1 | 31.2 9.2 | 35.2 | 60.0 | |
| TRIANGLE [5] | - | - | 68.7 4.2 | - | - | 59.9 3.7 | - | - | 21.9 5.0 | – | 50.2 | |
@@ 5.2 Affective Computing Benchmarks
We next evaluate on the MOSI, UR-FUNNY, and MUStARD datasets, which involve text, audio, and video for multimodal affect understanding. We assess (i) 11 retrieval, (ii) 21 retrieval, and (iii) multimodal classification via a linear probe. Table 2 shows that ConFu consistently improves cross-modal retrieval performance. In particular, the 21 retrieval setting surpasses 11 retrieval on UR-FUNNY and MUStARD, suggesting that combining two input modalities provides richer and more discriminative cues.
Notably, GRAM, the only baseline supporting both retrieval configurations, struggles with 11 alignments in general and fails for modalities M2 and M3 on MOSI and UR-FUNNY, suggesting that balancing 21 and 11 alignment is non-trivial. Furthermore, methods such as TRIANGLE and Symile, which focus on trimodal alignment without explicitly modeling pairwise relations, perform consistently worse than ConFu. Overall, ConFu ranks first or second across nearly all querytarget retrieval settings, in both 11 and 21 configurations, demonstrating robust and consistent cross-modal alignment. Moreover, Table 3 reports linear-probe classification results, showing that ConFu attains the highest accuracy on UR-FUNNY and remains competitive on MUStARD and MOSI.
@@ 5.3 Fine-Grained Bird Classification
ConFu is further evaluated on fine-grained bird classification using the VB100 and SSW60 datasets. Models are pretrained on Bird-MML, our curated dataset of 150K audio–image–text triplets and assessed under zero-shot and few-shot regimes. We consider two evaluation settings: (1) the multi-frame protocol from [30], where each video clip is divided into eight segments, the middle frame of each segment is sampled, and the mean embedding of the eight frames is used as the visual representation; and (2) a single-frame setting, providing a more constrained visual input per video. Table 4 reports zero-shot results in the multi-frame setting. ConFu achieves the highest accuracy on SSW60 (71.4%), confirming that multimodal fusion improves fine-grained discrimination when both modalities are informative. On VB100, where the audio modality is largely uninformative ( unimodal accuracy), the fused model (A+V) maintains performance close to the best visual-only baseline (18.15% vs. 20.69%), demonstrating robustness to distracting modalities. In contrast, GRAM, Symile, and TRIANGLE exhibit a notable decline in accuracy (5–10%) relative to ConFu. Few-shot trends in Fig. 4 mirror these findings. Fusion offers clear advantages on SSW60 and stable performance on VB100, showing that ConFu generalizes well even when one modality is weak.
@@
| Symile [27] | - | - | 61.4 0.2 | - | - | 13.2 0.4 |
| TRIANGLE [5] | - | - | 58.6 3.3 | - | - | 8.7 0.4 |
| GRAM [6] | 0.7 | 59.0 3.3 | 54.9 3.2 | 1.3 | 12.9 0.4 | 7.6 0.3 |
@@
In the single-frame setting, designed to evaluate performance with limited visual information, we averaged classification results across 19 uniformly sampled frames to reduce bias. As provided in Table 5, the fused (A+V) model exceeds the best baseline by roughly 4% on SSW60 and complementary audio–visual cues when visual input is sparse. On VB100, performance drops slightly (by 2–3%) due to the weaker audio modality ( accuracy). Overall, ConFu benefits from richer multimodal supervision, maintaining strong robustness under modality degradation while preserving scalability and simplicity. Additional results on a few-shot adaptation are in the Appendix C.1.
@@ 6 Ablation
Modality competition.
Although multimodal learning seeks to combine complementary cues, models often show modality dominance, where stronger signals suppress weaker ones [14]. To assess how our fusion is affected by this effect, we analyze in Fig. 5 the overlap of prediction errors across modalities and fusion. Results reveal visual dominance: The larger portion of correct samples (39%) are predicted by both the fusion and vision models, while the datapoints that only audio predicts correctly or fusion-audio overlap remain small (5–6%), and purely audio-driven successes are rare (2%). Nevertheless, about 5% of samples are uniquely solved by the audiovisual model, indicating that fusion captures complementary information beyond either modality alone. Finally, some classes benefit substantially from fusion (e.g., Label 31 see Appendix B.5), while others remain uniformly difficult.
Ablation on Noise-Induced Distribution Shift.
@@ 7 Discussion
Empirical results across synthetic and real-world datasets demonstrate that ConFu effectively captures complementary cross-modal structure. On the synthetic XOR benchmark, ConFu successfully models the underlying multimodal dependencies that pairwise and geometric baselines miss even with higher-dimensional representations (Appendix B.2). In AV-MNIST, ConFu dominates unimodal approaches in zero-shot classification, indicating successful utilization of both modalities. The 21 retrieval evaluations on UR-FUNNY, and MUStARD further confirm that combining two query modalities provides richer semantic grounding than either considered in isolation. An interesting observation in the 21 case on the MOSI dataset is that fusion slightly underperforms compared to the pairwise baseline. Remarkably, this issue is substantially alleviated by introducing partial modality masking (Appendix B.3), supporting the hypothesis that suboptimal performance stems from shortcut learning in cases where redundant information dominates.
@@
Limitations and Future Work. ConFu is well-positioned to scale to any number of modalities, offering flexibility for diverse multimodal tasks. However, as the number of modalities increases, computational demands may also rise. Depending on the task, selectively pruning certain alignment terms, shown in Appendix E, can help maintain efficiency. For example, omitting many-to-many alignments, when assuming that only single-modality queries or targets will be needed, opens a promising avenue for future adaptation and research.
ConFu currently relies on fully aligned modalities during training, and exploring ways to relax or bypass this requirement remains crucial for many applications. In particular, enabling ConFu to handle partially missing modalities, such as training with incomplete triplets or drawing on ideas related to pseudopair construction [35, 46, 36], could potentially make the framework more robust and broadly applicable across real-world multimodal settings. Further prospective directions could include extending ConFu to incorporate factorized representations inspired by recent factorization-based and redundancy-reduction approaches [19, 33, 8], developing adaptive mechanisms to regulate modality competition [18, 17, 24, 13, 38], and investigating how ConFu can enhance robustness in multimodal retrieval, particularly when one or more modalities are noisy, incomplete, or severely corrupted.
@@ 8 Conclusion
9 Acknowledgments
This work was supported by the HARSH project (project no. Y3TA - 0560901) within the framework of the National Recovery and Resilience Plan–Greece 2.0–with funding from the European Union–NextGenerationEU; the Hellenic Foundation for Research and Innovation (H.F.R.I.) under the “3rd Call for H.F.R.I.’s Research Projects to Support Faculty Members & Researchers” (H.F.R.I. Project Number: 26302); by the TITAN ERA Chair project (contract no. 101086741) within the Horizon Europe Framework Program of the European Commission; the METHUSALEM program (METH/26/003, “Methusalem-BioMedAI”); the Flemish Government AI Research Program (Leuven.AI – KU Leuven); the FWO project “Task- and Device-Agnostic Artificial Intelligence for EEG Analysis” (G046925N); and the Horizon Europe project “AI-PROGNOSIS” (Grant Agreement No. 101080581).
@@ A.2 MOSI, MUStARD, and UR-FUNNY
For these datasets, all models were trained for 100 epochs, producing 256-dimensional representations. We used a batch size of 128 and a learning rate of . Following prior work [19], each modality encoder was implemented as a 5-layer Transformer with 5 attention heads.
@@ B.1 Effect of the Parameter
As shown in Table 7, performance is not overly sensitive to . Accuracy remains stable while recall varies smoothly, indicating a controlled trade-off with minimal tuning required. Identifying a principled way to set or eliminate is an important direction.
| Dataset | Acc. (%) | |||
| MOSI | 0.2 | 67.7 1.7 | 19.5 1.0 | 20.1 1.9 |
| 0.5 | 66.7 2.1 | 19.5 1.5 | 19.6 2.0 | |
| 0.8 | 65.8 2.0 | 13.5 2.2 | 15.6 1.6 | |
| UR-Funny | 0.2 | 64.2 0.7 | 8.0 0.4 | 14.0 0.9 |
| 0.5 | 64.9 1.0 | 7.4 0.2 | 13.6 0.6 | |
| 0.8 | 64.6 0.7 | 6.9 0.3 | 13.2 0.7 | |
| MUSTARD | 0.2 | 59.7 3.2 | 42.1 2.2 | 57.1 2.5 |
| 0.5 | 64.1 2.5 | 45.3 2.6 | 62.6 3.1 | |
| 0.8 | 61.1 3.4 | 41.0 1.0 | 59.7 1.1 |
Regarding optimal values the results in Table 8 highlight the influence of the parameter, which balances the contribution of modality-specific and shared representations. These values were selected from the Pareto front obtained by jointly optimizing the two retrieval objectives - mean Recall@10 for 11 and 21 tasks. We observe that moderate values (0.1–0.5) yield the best trade-off between the two objectives, and that the optimal varies across datasets, indicating optimal depends on data distribution.
@@ B.4 Noise experiments
| Symile [27] | 58.4 | 21.2 | 59.5 | 27.9 | 60.0 | 40.9 |
| GRAM [6] | 58.9 | 4.0 | 58.3 | 8.94 | 58.1 | 25.9 |
| TRIANGLE [5] | 63.9 | 3.4 | 64.0 | 7.24 | 64.0 | 26.5 |
@@ C.1 Few shot adaptation on SSW60, VB100, CUB200 complete results
In this section, we present the complete results of our few-shot adaptation experiments on the SSW60, VB100, and CUB200 [32] datasets. The following figures include all baseline methods for comparison and illustrate adaptation performance across varying number of shot settings. These results provide a comprehensive view of how each method generalizes to limited-data scenarios and adapts to novel classes.
@@ D.2 Affective Computing Benchmarks
MultiBench [20] is a collection of benchmarking datasets designed to evaluate multimodal representation learning across a wide range of modalities and domains. In our experiments, we used the MOSI, UR-FUNNY, and MUStARD datasets. MultiBench provides pre-extracted features and predefined dataset splits, both of which we adopt in our setup.